OpenStudy (anonymous):

Find the first 5 terms of the sequence: an = (-1/2)n - 1

OpenStudy (anonymous):

Do you understand? Yes. No. Maybe.

OpenStudy (anonymous):

Maybe @some_someone

OpenStudy (anonymous):

Have you done the other ones like how i did the first one? Try it and see :)

OpenStudy (anonymous):

Okay. I'll try them @some_someone

OpenStudy (anonymous):

Hold on... I did it wrong. Lemme try again

OpenStudy (anonymous):

-1/2, 1/4, -1/8, 1/16 <----- This is my answer

OpenStudy (anonymous):

So find: when n = 1 when n = 2 when n = 3 when n = 4 when n = 5

OpenStudy (anonymous):

\[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\] when n = 1 \[a _{1} = \left( \frac{ -1 }{ 2 } \right)1 - 1\] \[a _{1} = \left( \frac{ -1 }{ 2 } \right)0\] \[a _{1} = 0 \]

OpenStudy (anonymous):

\[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\] when n = 2 \[a _{2} = \left( \frac{ -1 }{ 2 } \right)2 - 1\] \[a _{2} = \left( \frac{ -1 }{ 2 } \right)1\] \[a _{2} = \frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

I got my answer

OpenStudy (anonymous):

\[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\] when n = 3 \[a _{3} = \left( \frac{ -1 }{ 2 } \right)3 - 1\] \[a _{3} = \left( \frac{ -1 }{ 2 } \right)2\] \[a _{3} = -1\]

OpenStudy (anonymous):

what did you get @Emilyh117

OpenStudy (anonymous):

I got -1/2, 1/4, -1/8, 1/16 @some_someone

OpenStudy (anonymous):

\[0, -\frac{ 1 }{ 2 }, -1....\] these were the first three terms i found... You still need to find the other two. I didn't get what you got 0:

OpenStudy (anonymous):

Did you look at the first three i did?

OpenStudy (anonymous):

Yeah. Those weren't choices @some_someone

OpenStudy (anonymous):

hmmm were the ones you found choices?

OpenStudy (anonymous):

Yeah @some_someone

OpenStudy (anonymous):

Is this the given: \[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\]

OpenStudy (anonymous):

Yes @some_someone

OpenStudy (anonymous):

\[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\] when n = 4 \[a _{4} = \left( \frac{ -1 }{ 2 } \right)4 - 1\] \[a _{4} = \left( \frac{ -1 }{ 2 } \right)3\] \[a _{4} = -\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

\[a _{n} = \left( \frac{ -1 }{ 2 } \right)n - 1\] when n = 5 \[a _{5} = \left( \frac{ -1 }{ 2 } \right)5 - 1\] \[a _{5} = \left( \frac{ -1 }{ 2 } \right)4\] \[a _{5} = -2\]